OpenStudy (cookielate):

write each expression as the product of two factors : 2h+(6=h)+6*2 @sleepyjess

OpenStudy (anonymous):

I don't understand why the equals in brackets.

OpenStudy (cookielate):

oh sorry it supposed to be 2h+(6+h)+6*2

Directrix (directrix):

2h+(6+h)+6*2 simplifies to 3h + 18. So, factor out the 3: 3h + 18 = 3*(? + ?) @cookielate

OpenStudy (cookielate):

ok thanks

Directrix (directrix):

So, what did you get for an answer? @cookielate

OpenStudy (cookielate):

i got 24

Directrix (directrix):

No. 3h + 18 = 3*(h + 6) @cookielate

OpenStudy (cookielate):

oh okay thank you so much

Directrix (directrix):

You are welcome.